OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the points (9, 7) and (4, –8). y − = (x − 9) Complete the equation for the linear function whose graph contains the point (7, 8) and has a slope of −1.2. y = x + will fan and medal

OpenStudy (igreen):

First, we find the slope. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-8-7}{4-9}\) Simplify that

OpenStudy (anonymous):

\[\frac{ -15 }{ -5} = 3\]

OpenStudy (igreen):

Yes, so the slope is 3.. \(y - y_1 = m(x - x_1)\) Where y1 is the y-value, 'm' is the slope, and x1 is the x-value. We are already given the x-value, which is 9. And one of our points has x for 9..(9, 7). So our y-value is 7. So what should our equation be?

OpenStudy (anonymous):

y-7=3(x-9)

OpenStudy (igreen):

Yes! You got it

OpenStudy (anonymous):

can you also help with the last one? @iGreen

OpenStudy (igreen):

Sure, just close this and open a new one.